How much clay do you need to add to 6 ft^3 of soil with 30% clay in order to make a soil with 72% clay?
I suck at setting problems up, but can solve them fine. Having troubles with this one.
Thanks,
Tyler

Found 2 solutions by jorel1380, amalm06:
Answer by jorel1380(3719)

x+6(.3)=.72(x+6)
x+1.8=.72x+4.32
.28x=2.52
x=2.52/.28=9 cu.ft. of 100% clay to get the desired mix..

Answer by amalm06(224)

Use the method of alligation.
Let S denote the soil and C denote the clay.
Then 100-72=28 and 72-30=42
Therefore, the ratio of soil to clay is 28/42
Since there are 6 ft^3 of soil, (42/28)(6 ft^3) = 9 ft^3 (Answer)